3) What is the IRR for an investment that has a $5000 investment with $1000/year for 6 years?

PLEASE WRITE NEATLY AND EXPLAIN YOUR METHOD/SHOW ALL STEPS. THANKS!
Answer #1

Let i% be IRR, then

1000 * (P/A,i%,6) = 5000

(P/A,i%,6) = 5000 / 1000 = 5

using trail and error method

when i = 5%, value of (P/A,i%,6) = 5.075692

when i = 6%, value of (P/A,i%,6) = 4.917324

using interpolation

I = 5% + (5.075692 - 5) / (5.075692 - 4.917324)*(6% -5%)

I = 5% + 0.47%

i = 5.47% (Approx)